### v2

# ALOE: Action-Level Off-Policy Evaluation for Vision-Language-Action Model Post-Training

Abstract
We study how to improve large foundation vision-language-action (VLA) systems through online reinforcement learning (RL) in real-world settings. Central to this process is the value function, which provides learning signals to guide VLA learning from experience. In practice, the value function is estimated from trajectory fragments collected from different data sources, including historical policies and intermittent human interventions. Estimating the value function of current behavior quality from the mixture data is inherently an off-policy evaluation problem. However, prior work often adopts conservative on-policy estimation for stability, which avoids direct evaluation of the current high-capacity policy and limits learning effectiveness. In this paper, we propose ALOE, an action-level off-policy evaluation framework for VLA post-training. ALOE applies chunking-based temporal-difference bootstrapping to evaluate individual action sequences instead of predicting final task outcomes. This design improves effective credit assignment to critical action chunks under sparse rewards and supports stable policy improvement. We evaluate our method on three real-world manipulation tasks, including smartphone packing as a high-precision task, laundry folding as a long-horizon deformable-object task, and bimanual pick-and-place involving multi-object perception. Across all tasks, ALOE improves learning efficiency without compromising execution speed, showing that off-policy RL can be reintroduced in a reliable manner for real-world VLA post-training. Videos and additional materials are available at our project website https://rooshy-yang.github.io/aloe.
Figure 1: Overview of our real-world actor–critic framework for VLA post-training. Left: Our method adopts an actor–critic framework in which the actor is a flow-matching–based foundation VLA model and the critic is a lightweight ensemble Q-network. The actor outputs action sequences for online real-world rollouts, while the critic predicts ensemble Q-values to assess the quality of the actor’s action chunks under the current observation. Middle: Real-world RL is conducted in three stages. (1) Data collection under human intervention: the VLA policy is first warm-started with offline behavior cloning and then deployed on real robots. Both successful and failed rollouts are stored in buffer. When failures or unsafe behaviors occur, a human intervenes via teleoperation to take over from the failure state and guide the robot to the goal state. (2) Off-policy critic estimation: the critic is trained on the aggregated dataset using Q-chunking TD updates. (3) Policy improvement: the actor is optimized using pessimistic value estimation and advantage-weighted maximum likelihood. Right: We evaluate the method on real-world manipulation tasks and demonstrate improvements in task success rate, robustness to disturbances, and zero-shot generalization to unseen objects.
I Introduction
Vision-language-action (VLA) models have emerged as a promising paradigm for robotic manipulation, enabling policies to condition actions on both real-world visual observations and natural language instructions [1, 5, 4, 52, 57, 26, 44, 25]. Recent progress has shown that combining imitation learning with reinforcement learning (RL) can substantially improve task performance beyond supervised learning [14, 51, 15, 12]. A key factor behind these improvements is the value function, which evaluates action quality and provides optimization signals for VLA policies. Value functions play a central role in RL by reducing variance and improving sample efficiency [42, 43]. However, accurately estimating value functions in real-world VLA systems remains particularly challenging [14, 32]. Unlike training on simulation, real-world RL for VLA systems faces practical constraints such as data scarcity, partial observation on real-world dynamics, and limited environment parallelism. These factors make large-scale data collection costly and slow down RL iteration for policy convergence. Moreover, real-world VLA datasets typically consist of a heterogeneous mixture of sources, including rollouts from online deployment, corrective interventions from human operators in response to errors or unsafe behaviors, and historical data from earlier policies or pre-collected demonstrations. In such settings, value estimation for current VLA policy is inherently an off-policy problem, but it can suffer from error accumulation due to distributional uncertainty [19, 47], leading to biased estimates and unstable learning dynamics. To maintain stability for high-capacity VLA systems, prior work often adopts conservative learning strategies based on on-policy value estimation [14, 36]. Some compromise approaches such as AWR and RECAP [36, 14] relax strict on-policy requirements. In these methods, the learned value function is not intended to directly evaluate the current policy being optimized. Instead, it estimates the expected return of an implicit policy induced by training data from multiple sources. As a result, the value function may not correspond to any single executable policy, but rather reflects the average behavior distribution on which it is trained and is used to guide current policy behavior.
While this on-policy estimation design improves training stability, its application to real-world VLA systems relies on a strong assumption: the behavior distribution of the evolving policy remains close to the mixture distribution represented in the replay buffer. This assumption is particularly fragile in real-world VLA settings where human interventions, policy updates, and environment variability continuously shift the data distribution. Under this assumption, the value function trained on historical data can serve as an approximate proxy for evaluating the current policy’s actions. However, when the VLA policy evolves and its behavior distribution deviates substantially from that mixture, the learned value function may no longer reflect the true value of the current policy. This leads to misleading optimization signals and potentially incorrect training directions. For example, when a policy learns new recovery behaviors that were absent in historical data, the value function trained on past trajectories may underestimate these actions. This discourages exploration of corrective behaviors that are essential for robust real-world manipulation. This compromise induces a fundamental trade-off between robustness and learning efficiency in real-world VLA training, and motivating the following question:
When can off-policy reinforcement learning be reliably applied to improve real-world VLA policies?
In this work, we introduce ALOE, an action-level off-policy evaluation framework that enables effective off-policy policy improvement under real-world RL setting. We show that such conservatism from prior approaches is not inherent to real-world VLA training, but instead arises from specific assumptions about how value estimates are obtained and used. By explicitly addressing key sources of critic bias, the assumptions can be partially relaxed without compromising training stability. Specifically, we reintroduce traditional temporal-difference bootstrapping method to estimate the behavior quality of current policy, providing effective off-policy policy evaluation and stitch different fragments from different source. To accelerate credit propagation under long-horizon and sparse reward setting, we adopt Q-chunking for efficient reward signal backpropagation, ensuring efficient propagation of reward signals to critical action frames. To stabilize off-policy Q learning, we consider a pessimistic value estimation for uncertainty, mitigating the error accumulating on uncertain behavior. We then adopt this value function as preference signal to guide high-capacity policy update, providing stable policy improvement. Our approach allows VLA policies to leverage fine-grained, action-level value information while preserving the robustness properties essential for real-world training. We summarize our contributions as follows:
  • We propose an off-policy policy evaluation framework for real-world VLA training under human-in-the-loop data collection, enabling action-level evaluation to support reliable policy improvement.
  • We develop stabilization mechanisms for action-value learning in real-world settings, enabling reliable long-horizon credit assignment for flow-based VLA policies.
  • We evaluate our approach on three real-world robotic manipulation tasks, demonstrating improved task success, generalization, learning efficiency, and recovery from execution errors compared to prior VLA training methods.
II Related Work
II-A Foundations of Policy Evaluation
Policy evaluation aims to estimate the expected long-term outcomes of actions or policies and is commonly categorized along two axes: on-policy vs. off-policy estimation, and Monte Carlo (MC) targets vs. Temporal Difference (TD) bootstrapping targets [42]. On-policy methods estimate the expected return of the action distribution induced by the data-collecting policy, ensuring alignment with the data distribution and stable learning [37, 48, 22]. In contrast, off-policy methods estimate the expected return of specific actions or target policies using data collected from different behavior policies, enabling broader action coverage and improved sample efficiency [21, 16, 7, 55, 56]. MC-based methods rely on complete episodic returns and suffer from high variance, while TD-based methods leverage Bellman bootstrapping to learn from fragment trajectories [42]. Although off-policy, TD-based value estimation is well studied in classical RL, extending these mechanisms to high-capacity VLA models in real-world settings remains challenging [14, 47, 19]. Our work mitigates these challenges by enabling reliable action-value estimation for large diffusion- and flow-based VLA policies on more complex real-world tasks.
II-B Reinforcement Learning for Large VLA Policies
Recent work has explored scaling reinforcement learning to large VLA models on real robotic systems [45, 29, 27, 6, 23]. Many approaches adopt on-policy optimization, such as PPO or REINFORCE [49, 38, 48], prioritizing stability but limiting value estimation to the current behavior distribution. Another prevalent line of work employs trajectory-level preference or progress modeling, where a state-value function predicts task completion progress and serves as a reweighting signal for imitation-style updates [14, 53, 10, 33, 31]. While effective for stabilizing training, these methods do not explicitly evaluate individual actions and provide limited action-level guidance. Action-value and advantage-based methods for VLA have also been studied [13, 8, 18], but are largely restricted to offline RL settings with static data. In contrast, real-world VLA training typically follows an iterated RL paradigm with continuous online interaction and human-in-the-loop interventions [41, 35, 30, 20]. Existing methods in this regime generally avoid off-policy action-value estimation due to stability concerns. Our work bridges this gap by introducing a controlled, off-policy, action-level value estimation framework for more complex tasks like long-horizon clothing folding, high-precision phone packing tasks, enabling fine-grained action correction and trajectory stitching while preserving robustness under online, human-in-the-loop deployment.
III Preliminaries
III-A Vision-Language-Action Model as an Episodic MDP
We model VLA as an episodic Markov Decision Process (MDP) [42] defined by the tuple $(\mathcal{S},\mathcal{A},p,r,\gamma)$. At each timestep $t$, the agent observes a state $s_{t}\in\mathcal{S}$, which includes visual observations and proprioceptive information, and receives a natural language instruction $\ell$ specifying the task. The instruction $\ell$ is treated as a conditioning variable of VLA. The agent selects an single action $a_{t}\in\mathcal{A}$ according to a VLA policy $\pi_{\theta}(a_{t}\mid s_{t},\ell)$ parameterized by $\theta$. The environment transitions according to unknown dynamics $p(s_{t+1}|s_{t},a_{t})$ and emits a reward $r_{t}=r(s_{t},a_{t})$. We consider episodic tasks that terminate upon either task success or failure, and denote termination by $d_{t}\in\{0,1\}$. The discount factor is $\gamma\in(0,1)$. The objective of reinforcement learning is to learn a policy that maximizes the expected discounted return of the current policy,
$J(\pi_{\theta})=\mathbb{E}_{\tau\sim\mathcal{D},\pi_{\theta}}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right],$ (1)
where trajectories $\tau$ are sampled from a replay buffer $\mathcal{D}$ containing off-policy transitions $(s_{t},a_{t},r_{t},s_{t+1},d_{t},\ell)$ collected by previous policies. While this objective can in principle be optimized using on-policy policy gradient methods [48, 40, 37], real-world VLA training is often constrained by costly environment parallelism, long-horizon task structure, and sparse rewards, which motivate the efficient reuse of history data.
III-B Actor-Critic Framework
Actor-critic methods [11, 43] address these challenges by maintaining two components: an actor $\pi_{\theta}$ that represents the VLA policy, and a critic that estimates long-term outcomes to guide policy updates. We begin by defining the action-value function of a policy $\pi$. For a given state-action pair $(s,a)$ and language instruction $\ell$, the action-value function is defined as
$Q^{\pi}(s,a,\ell)=\mathbb{E}_{\pi}\!\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\;\middle|\;s_{0}=s,\;a_{0}=a,\;\ell\right],$ (2)
which represents the expected discounted return obtained by first taking action $a$ at state $s$ under instruction $\ell$, and then following policy $\pi$ thereafter. The state-value function is defined as the expected action-value under the policy,
$V^{\pi}(s,\ell)=\mathbb{E}_{a\sim\pi(\cdot\mid s,\ell)}\!\left[Q^{\pi}(s,a,\ell)\right],$ (3)
and summarizes the expected return of state $s$ under instruction $\ell$ when actions are sampled from $\pi$. Given an action-value function $Q(s,a,\ell)$ and a state-dependent baseline $V(s,\ell)$, the advantage function is defined as
$A^{\pi}(s,a,\ell)=Q^{\pi}(s,a,\ell)-V(s,\ell),$ (4)
which measures the relative quality of an action compared to the policy’s average behavior at a given state under instruction $\ell$. The advantage function is commonly used to guide policy improvement by increasing the likelihood of actions with positive advantage while suppressing suboptimal actions. In practice, the critic is often learned from a replay buffer $\mathcal{D}$ containing transitions $(s,a,r,s^{\prime},d,\ell)$.
IV Method
Algorithm 1 ALOE: Action-Level Off-Policy Evaluation Framework for VLA Policies
Require: initial VLA policy $\pi_{\theta}$; critic ensemble $\{Q_{\phi_{i}}\}_{i=1}^{K}$ with targets $\{Q_{\phi_{i}^{-}}\}_{i=1}^{K}$; replay buffer $\mathcal{D}$; discount $\gamma$; hyperparameters $(\beta,\epsilon_{\mathrm{clip}},\omega)$; update steps $(N_{Q},N_{\pi})$.
1:Train $\pi_{\theta}$ by imitation learning on successful demonstrations.
2:for $k=1$ to $T$ do
3:  Rollout $\pi_{\theta}$ under instruction $\ell$ with human intervention and early termination; store all transitions into $\mathcal{D}$.
4:  for $j=1$ to $N_{Q}$ do
5:   Sample minibatch $\mathcal{B}\subset\mathcal{D}$ and corresponding next actions from $\pi_{\theta}$.
6:   Compute TD target $y$ using Eq. (7).
7:   Update critic ensemble by minimizing Eq. (8).
8:   Update target networks $\phi_{i}^{-}\leftarrow\omega\phi_{i}+(1-\omega)\phi_{i}^{-}$.
9:  end for
10:  for $j=1$ to $N_{\pi}$ do
11:   Sample minibatch $\mathcal{B}\subset\mathcal{D}$, and compute corresponding $Q_{\mathrm{pess}}(s_{t},\mathbf{a}_{t:t+h},\ell)$ using Eq. (9).
12:   Estimate $V(s,\ell)$ by sampling $a\sim\pi_{\theta}(\cdot\mid s,\ell)$ and compute advantage using Eq. (10).
13:   Compute weight $w$ using Eq. (11).
14:   Update policy $\pi_{\theta}$ by minimizing Eq. (IV-C).
15:  end for
16:end for
17:Return: trained policy $\pi_{\theta}$.
Our method targets real-world VLA reinforcement learning under human-in-the-loop data collection, where experience consists of fragmented and policy-mixed trajectories due to safety interventions. We decompose learning into policy evaluation and policy improvement. For policy evaluation in Sec. IV-A, we learn an off-policy action-value function using TD learning with Q-chunking and a pessimistic ensemble, enabling efficient long-horizon credit propagation and robust value estimation. For policy improvement in Sec. IV-B, we treat the critic as a relative preference signal and perform advantage-weighted policy updates. Finally, the detailed implementation in Sec. IV-C discusses how to adapt the RL framework into flow-based VLA architecture. The resulting algorithm enables reliable and data-efficient post-training of flow-based VLA policies in Algorithm 1.
Figure 2: Illustrations of the real-world manipulation tasks and robot setup used in our experiments. We evaluate on phone packing, laundry folding, and multi-object pick and place, showing representative task stages, object variations, and the physical robot platform with three RGB cameras.
IV-A Off-Policy Critic Estimation under Real-World Data Collection
We first describe the policy evaluation component of our method. As shown in Fig. 1, we follow prior real-world VLA training paradigms [14, 46, 7, 36, 51]. Data is incrementally collected through repeated rollouts of a continually evolving VLA policy. During online deployment, rollouts are frequently interrupted when the policy executes unsafe behaviors, such as collisions or joint limit violations. In such cases, the trajectory is truncated, labeled with a failure outcome, and a human operator intervenes to correct the robot’s actions before execution resumes. This process may repeat multiple times until the task is successfully completed. As a result, the replay buffer $\mathcal{D}$ contains not only data generated by a mixture of behavior policies, but also fragmented trajectories in which different segments of a single task execution may originate from different policies or human corrections. Consequently, learning from $\mathcal{D}$ is inherently off-policy with respect to the current actor, and long-horizon task executions are rarely observed as contiguous rollouts generated by a single policy.
Under this setting, our objective is to estimate the expected long-term outcome of taking an action at a given state under a language instruction, i.e., the action-value function in Eq. (2). A common alternative in real-world VLA systems is to use trajectory-level MC returns as a value signal [14, 22], which assign value by accumulating observed future rewards along the same trajectory and implicitly assume access to a complete trajectory suffix. In real-world training, this assumption is frequently violated due to safety-driven truncation and human intervention, resulting in fragmented task executions collected under heterogeneous behavior policies. Consequently, the future return following an action is often only partially observed, making it difficult for MC-based estimators to assign reliable credit under sparse rewards. This limitation is particularly severe in long-horizon tasks, where informative trajectory fragments must be combined across policies to reason about outcomes that are never observed as contiguous rollouts.
To mitigate these challenges, we adopt TD learning [42], which bootstraps from subsequent value estimates and allows value information to propagate backward through time. Starting from the Bellman equation:
$Q^{\pi}(s,a,\ell)=r+\gamma\,\mathbb{E}_{s^{\prime}\sim p,\,a^{\prime}\sim\pi}\left[Q^{\pi}(s^{\prime},a^{\prime},\ell)\right].$ (5)
TD learning replaces the expectation over future trajectories with a single-step target constructed from observed transitions. Given a transition $(s_{t},a_{t},r_{t},s_{t+1},d_{t},\ell)\sim\mathcal{D}$, we define the TD target
$y_{t}=r_{t}+\gamma(1-d_{t})\,Q_{\phi^{-}}(s_{t+1},a_{t+1},\ell),$ (6)
where $\phi^{-}$ denotes the parameters of a target network, and the next action is sampled from the current policy $a_{t+1}\sim\pi_{\theta}(\cdot\mid s_{t+1},\ell)$.
The advantage of TD learning over MC estimation in this setting arises from the contraction property of the Bellman operator [3]. Unlike MC estimation, which requires observing a complete trajectory $\tau={(s_{t},a_{t})}_{t=0}^{H}$ to assign value to an initial state–action pair, TD learning applies the Bellman update locally at each transition. This locality allows value information to propagate incrementally through the state–action space, even when trajectories are fragmented or truncated. As a result, the process allows value information from different fragments to be implicitly stitched together, allowing the critic to aggregate credit from disjoint real-world rollouts and assign meaningful long-horizon credit without requiring contiguous task executions. This alleviates reliance on manual trajectory organization and human annotation under real-world setting. Moreover, TD learning estimates an action-conditional value function $Q(s,a,\ell)$ instead of $V(s)$. This enables fine-grained, action-level credit assignment. Such action-level feedback is critical for identifying failure-prone decisions and key states in long-horizon manipulation tasks with complex intermediate dynamics.
While TD learning enables action-level credit assignment, standard one-step TD updates propagate value information backward by only a single timestep. In long-horizon manipulation tasks with sparse terminal rewards, this leads to slow credit propagation and inefficient learning, as meaningful reward signals may need to traverse about thousands of transitions. To address this issue, we adopt a Q-chunking [24] value backup that propagates value over multiple timesteps while remaining unbiased. Specifically, we define a chunked action sequence $\mathbf{a}_{t:t+h}\triangleq(a_{t},a_{t+1},\dots,a_{t+h-1}),$ and learn a chunked action-value function $Q(s_{t},\mathbf{a}_{t:t+h},\ell)$ that evaluates the return of executing this exact action sequence from state $s_{t}$ under instruction $\ell$. For a transition segment $(s_{t},\mathbf{a}_{t:t+h},r_{t:t+h-1},s_{t+h},\ell)\sim\mathcal{D}$, the Q-chunking Bellman target is defined as
$y_{t}^{(h)}=\sum_{k=0}^{h-1}\gamma^{k}r_{t+k}+\gamma^{h}\mathbb{E}_{\mathbf{a}^{\prime}\sim\pi_{\theta}(\cdot\mid s_{t+h},\ell)}\big[Q_{\phi^{-}}(s_{t+h},\mathbf{a}^{\prime},\ell)\big],$ (7)
where $\mathbf{a}^{\prime}=\mathbf{a}_{t+h:t+2h}$ denotes the next action chunk sampled from the current policy. Unlike standard $n$-step TD, the backup uses the exact action sequence that generated the intermediate rewards, eliminating the mismatch between the return estimator and the behavior policy. As a result, Q-chunking achieves faster value propagation while preserving unbiased estimation [24].
In regions of the state-action space with insufficient observations, we find that traditional TD-based methods may extrapolate aggressively and produce overestimated action values. This effect is amplified in safety-critical manipulation tasks, where exploration is constrained and data diversity is inherently limited. To mitigate the overestimation arising from uncertainty, we employ a pessimistic ensemble of $K$ action-value functions $\{Q_{\phi_{i}}\}_{i=1}^{K}$. Each critic is trained using the same chunked TD target in Eq. (7), by minimizing the objective:
$\mathcal{L}_{\mathrm{critic}}=\frac{1}{K}\sum_{i=1}^{K}\mathbb{E}_{\mathcal{D}}\left[\big(Q_{\phi_{i}}(s_{t},\mathbf{a}_{t:t+h},\ell)-y^{(h)}_{t}\big)^{2}\right],$ (8)
where $y^{(h)}_{t}$ denotes the chunked TD target. At policy update time, we form a conservative estimate:
$Q_{\mathrm{pess}}(s_{t},\mathbf{a}_{t:t+h},\ell)=\min_{i}Q_{\phi_{i}}(s_{t},\mathbf{a}_{t:t+h},\ell),$ (9)
which approximates a lower confidence bound on the true action value [9, 28]. This aggregation evaluates the current policy while providing more sensitive identification of uncertainty, improving robustness under real-world deployment.
IV-B Advantage-Weighted Policy Improvement for VLA Policies
We now describe how the learned action-value function is used to update the VLA policy. Although the critic is learned off-policy, our value evaluation integrates naturally with the constrained policy update strategy used in RECAP [14], treating the critic as a relative preference signal. This preserves constraint updates and ensures stable policy improvement across online iterations.
Figure 3: Visualization of Q-value. Q-values predicted by the learned off-policy critic along a representative trajectory from the target policy (50Hz). The critic assigns sharply lower values to actions leading to failure and higher values to successful behaviors, demonstrating fine-grained action-level credit assignment in long-horizon manipulation tasks.
Given a transition, we evaluate the quality of the data action relative to the current policy $\pi$ by defining the advantage:
$A^{\pi}(s_{t},\mathbf{a}_{t:t+h},\ell)=Q_{\mathrm{pess}}(s_{t},\mathbf{a}_{t:t+h},\ell)-\mathrm{sg}(V^{\pi}(s_{t},\ell)),$ (10)
where $V^{\pi}(s_{t},\ell)=\mathbb{E}_{\mathbf{a}_{t:t+h}^{\prime}\sim\pi_{\theta}(\cdot\mid s_{t},\ell)}[Q_{\mathrm{pess}}(s_{t},\mathbf{a}_{t:t+h}^{\prime},\ell)]$ denotes the state-value function of the current policy. The stop-gradient operator $\mathrm{sg}$ prevents gradients from flowing through the last term, decoupling policy optimization from critic learning and improving training stability. By construction, this advantage provides a local, policy-dependent improvement signal that compares the action observed in the data against the behavior of the current policy at the same state and instruction. Policy improvement is performed via an advantage-weighted log-likelihood objective. To emphasize relative action quality while preserving stability, we transform the advantage into a non-negative weight using a clipped exponential,
$w(s_{t},\mathbf{a}_{t:t+h},\ell)=\exp\!\left(\mathrm{clip}\!\left(\frac{A^{\pi}(s_{t},\mathbf{a}_{t:t+h},\ell)}{\beta},-\epsilon_{\mathrm{clip}},\epsilon_{\mathrm{clip}}\right)\right),$ (11)
where $\beta$ controls the sharpness of the weighting and $\epsilon_{\mathrm{clip}}$ limits the maximum influence of any single transition, analogous to trust-region clipping in PPO-style updates [48]. The policy is then updated by maximizing the weighted log-likelihood over the entire replay buffer,
$\mathcal{L}_{\mathrm{actor}}=\mathbb{E}_{\mathcal{D}}\left[w(s_{t},\mathbf{a}_{t:t+h},\ell)\,\log\pi_{\theta}(\mathbf{a}_{t:t+h}\mid s_{t},\ell)\right].$ (12)
This update rule has several desirable properties. First, policy improvement is explicitly regularized. The policy is encouraged to increase the likelihood of actions observed in the data, which prevents uncontrolled extrapolation beyond the replay buffer. Second, because the advantage is defined relative to the current policy, the update focuses on correcting local deficiencies of the policy while remaining robust to global value estimation errors. These properties suggest that the update in Eq. (12) implements a conservative but principled policy improvement step:
Theorem IV.1 (Constraint Policy Improvement).
For a given state $s$, let $\pi_{\mathrm{ref}}$ be the implicit behavior policy representing the data distribution in $\mathcal{D}$. The optimal solution to the advantage-weighted objective is the analytical solution to the following constrained optimization problem:
$\max_{\pi}$ $\mathbb{E}_{\mathbf{a}_{t:t+h}\sim\pi(\cdot\mid s_{t},\ell)}\big[Q(s_{t},\mathbf{a}_{t:t+h},\ell)\big]$ (13)
s.t. $D_{\mathrm{KL}}\!\left(\pi(\cdot\mid s_{t},\ell)\,\|\,\pi_{\mathrm{ref}}(\cdot\mid s_{t},\ell)\right)\leq\epsilon.$
The constraint $\epsilon$ is implicitly controlled by the temperature parameter $\beta$.
Theorem IV.1 formalizes this interpretation by showing that the resulting policy update corresponds to the optimal solution of a constrained policy optimization problem, where improvement is explicitly restricted to the support of the data distribution.
Figure 4: The average success rate of the three manipulation tasks. We evaluate the baselines under real-world setting with several runs and calculate the average success rate.
Figure 5: Evaluation of efficiency, zero-shot generalization, and robustness across tasks. The left plot reports execution throughput on the phone-case task, where higher throughput indicates more efficient task completion. The middle plot evaluates generalization to unseen objects in a pick-and-place setting. The right plot measures robustness by injecting disturbances into the robot’s actions and assessing its ability to recover and successfully complete the task.
IV-C Implementation Details
We adopt the $\pi_{0.5}$ flow-matching VLA model [4] as the actor. Given the visual observation, robot proprioception and language instruction, the actor predicts continuous action chunks via the flow-matching action head. We fine-tune the actor end-to-end during RL. For critic, we adopt $\pi_{0.5}$ pretrained SigLIP visual encoder [54] to extract image features, and fuse with other modalities using a Transformer backbone. To improve stability and reduce compute, we freeze the SigLIP encoder. We use an ensemble of $K$ Q-functions implemented with $K$ learnable Q readout tokens appended to the Transformer input sequence. Each readout token is mapped by an independent Q-head to predict a scalar Q value. We adopt a sparse terminal reward with a per-step penalty, following prior VLA RL practice [14]. For an episode terminating at step $T$, the reward is
$r_{t}=\begin{cases}0,&t=T\ \text{and success}\\ -C_{\text{fail}},&t=T\ \text{and failure}\\ -1,&\text{otherwise},\end{cases}$ (14)
where we set $C_{\text{fail}}$ to the average episode length according to the tasks.
Our actor parameterizes continuous action chunks via flow matching, so the continuous log-likelihood term in Eq. (12), $\log\pi_{\theta}(\mathbf{a}_{t:t+h}\mid s_{t},\ell)$, is not tractable to evaluate exactly. Following prior flow/diffusion-style VLA training practice [14], we optimize a tractable surrogate by treating the negative flow-matching objective as a proxy for the continuous log-likelihood, i.e.,
$\mathcal{L}_{\mathrm{actor}}$ $=\mathbb{E}_{\mathcal{D}}\Big[w(s_{t},\mathbf{a}_{t:t+h},\ell)\bigl\lVert\boldsymbol{\epsilon}-\mathbf{a}_{t:t+h}-f_{\theta}\!\bigl(\tilde{\mathbf{a}}_{t:t+h},s_{t},\ell\bigr)\bigr\rVert_{2}^{2}\Big],$
$(s_{t},\mathbf{a}_{t:t+h},\ell)\sim\mathcal{D},\quad\eta\sim\mathcal{U}[0,1],\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (15)
where the noised action defined as $\tilde{\mathbf{a}}_{t:t+h}=\eta\mathbf{a}_{t:t+h}+(1-\eta)\boldsymbol{\epsilon}$. Intuitively, maximizing the weighted policy log likelihood corresponds to minimizing weighted flow-matching loss for the action chunks under the same inputs. The state input to the actor and critic is $s_{t}=[X^{1}_{t},X^{2}_{t},X^{3}_{t},q_{t}]$, where $X^{i}_{t}$ denotes the image from the $i$-th camera and $q_{t}$ denotes the robot proprioceptive observation. More training details and parameter settings can be found on Appendix -C.
V Experiments
In this section, we evaluate ALOE on a set of real-world, robotic manipulation tasks designed to stress long-horizon reasoning, precise action selection, and robustness under human-in-the-loop data collection. Specifically, we consider three representative tasks: Pack Smart Phone, Folding Laundry, and Product Sorting. These tasks require a combination of fine-grained manipulation, multi-stage task composition (e.g., pick–place–push), and reasoning over visually ambiguous intermediate states. All experiments are conducted on a real robotic platform, illustrated in Fig. 2. More experimental implementation can be found on Appendix -B.
V-A Description of Tasks
Pack Smart Phone
The Pack Smart Phone task requires the robot to align and attach a phone case onto a device body. Successful completion involves precise pose alignment, careful force control during insertion, and recovery from intermediate misalignment. The task is characterized by sparse rewards: the episode is marked successful only if the phone case is fully attached without collision-induced failure. Early termination is triggered when unsafe contact or irreversible misalignment is detected, often followed by human intervention. This task is particularly challenging for trajectory-level value estimation, as failure is often caused by a single suboptimal action taken during a narrow time window.
Folding Laundry
Folding Laundry is a long-horizon manipulation task involving deformable object dynamics. The robot must grasp, lift, reposition, and fold a piece of cloth into a target configuration. This task emphasizes the need for precise grasp point selection, continuous visual feedback, and long-horizon planning to handle the highly variable shape and state of the cloth.
Product Sorting
The Product Sorting task requires the robot to identify, manipulate, and place multiple objects with diverse shapes and visual appearances into designated bins. The task involves object recognition, sequential decision-making, and collision-aware motion planning. Episodes may terminate early due to object drops or workspace violations. This setting evaluates the learning algorithm’s ability to generalize from fragmented experience and to assign credit to individual actions that influence downstream task success.
Figure 6: The performance of each RL iteration. The success rates of ALOE on the Phone Case and Dual-Arm Object tasks increase consistently as RL training progresses, demonstrating stable policy improvement across online iterations.
V-B Comparisons and Ablations
We compare ALOE against the following baselines:
  • BC ($\pi_{0.5}$ [4]): A pure imitation learning baseline trained on successful trajectories using the $\pi_{0.5}$ architecture. This baseline reflects the performance achievable without reinforcement learning. The BC is used as warm up for all baselines.
  • DAgger [39]: Based on the $\pi_{0.5}$, an iterative imitation learning method that aggregates online rollouts and human corrections. While effective at reducing covariate shift, DAgger does not explicitly leverage value estimation for long-term credit assignment.
  • AWR [36]: An on-policy RL approach that uses Monte Carlo estimation to evaluate the reference policy in the replay buffer. AWR employs state-level value estimation and prioritizes robustness under real-world data collection but does not exploit action-level value information.
V-C Quantitative results
What Does the Q-Function Actually Learn
To better understand the role of the learned action-value function, We analyze the Q-value trend from a trajectory sampled from target policy in Fig. 3. For example, the color-coded regions on the left one correspond to behaviors critical for task success. The Q-value in the red region drops sharply when the robot fails to scoop up the phone, demonstrating the critic’s ability to identify failure modes. The Q-value in the green region increases sharply, indicating successful recovery behavior. The Q-value in the orange region increases gradually with fluctuations, assigning fine-grained credits to individual actions and providing richer supervision than trajectory-level value estimation.
How much does ALOE improve the policy
To evaluate the effectiveness of our off-policy algorithm, we compare task success rates across all methods, as summarized in Fig. 4. ALOE consistently outperforms all baselines on all three tasks, achieving higher success rates with fewer environment interactions compared to AWR, which demonstrates improved learning efficiency. For the Pack Smart Phone task, which requires high precision, we additionally measure throughput, which is the number of tasks completed per hour as shown in Fig. 5. Our method achieves higher throughput, indicating that the improved success rate translates to more efficient task execution in practice. For the Product Sorting task, we further evaluate generalization performance on unseen objects. ALOE maintains robust performance on novel objects, demonstrating better generalization capability compared to baselines.
How much does ALOE improve flow-based VLA over multiple iterations
We evaluate how different learning methods improve a flow-based VLA policy across multiple training iterations, as shown in Fig. 6. Consistent with prior work, AWR yields noticeable gains over pure imitation learning in early iterations by stabilizing policy updates under real-world data collection. However, its improvements tend to steady and saturate as training progresses, particularly in tasks with delayed or action-specific failure modes such as Folding Laundry. In contrast, ALOE continues to improve the same flow-based VLA backbone over successive iterations, achieving higher final success rates and better sample efficiency. This result indicates that action-level off-policy value estimation provides additional learning signal beyond trajectory-level preference modeling when refining flow-based VLA policies.
VI Conclusion
In this paper, we analyze how value functions can be more effectively leveraged in real-world VLA RL. We propose ALOE, an action-level off-policy value estimation method that enables policy improvement while preserving robustness. Empirically, we demonstrate that off-policy, Bellman-based action-value estimation can be made reliable under human-in-the-loop data collection, and that ALOE consistently improves learning efficiency and task throughput over prior conservative VLA training methods across three real-world manipulation tasks. In future work, we plan to explore combining powerful off-policy value estimation with automated task and environment reset pipelines to enable fully autonomous real-world VLA training.